# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

### v1

@@ 5.4 Results
1. Even advanced models cannot achieve success without any hints. Table 3 presents the results of various VLMs integrated within the VisEscaper framework under both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Without any hint($exp_{\text{base}}$), even Claude-3.5-Sonnet achieves only a 21.7% escape success rate and 32.37% goal completion, which is still 3-4 times lower than human scores. Considering that humans show minimal difference between achievement rate and final escape success rate (97.5 for GC and 86.0 for SR), the results from all models demonstrate that despite making some progress within each room, they cannot guarantee long-horizontal continuous success when no guidance is provided.
2. High-performing models show a drastic increase in all metrics, achieving objectives with fewer hints and within fewer turns. When hints are provided to stuck agents ($exp_{\text{hint}}$), there is a significant increase in success rate and goal completion across all models. Notably, Claude-3.5-Sonnet and GPT-4o escaped rooms using less than half the average steps compared to other models, making it much more efficient at path optimization, showing the greatest improvement especially in SPL among all metrics. Also, smaller models still improve substantially with hints, but rely on them more frequently, as shown by higher HCR.
3. High-performing models demonstrate accelerated goal completion rates in later stages compared to early stages. Figure 6 illustrates the progression of GC in Table 3 across all rooms during the first 100 steps. Notably, while Claude-3.5-Sonnet performs best overall, it exhibits a distinctly accelerated trajectory after approximately 30-40 steps. This acceleration indicates that strong models better manage accumulated room-state changes, effectively leveraging previous observations to make strategic and efficient decisions later in the trajectory. Additionally, models exhibiting rapid progress (Claude-3.5-Sonnet and GPT-4o) actually used fewer hints compared to other models(HCR of $exp_{\text{hint}}$ in Table 3), indicating that high-performing models do not simply rely on hints throughout escape-room tasks. but accelerate progress by effectively leveraging organized prior information after resolving specific obstacles.
Figure 6: Average GC(Goal Completion) progression across all rooms and experiments during the first 100 steps. X-axis represents the number of steps, and the Y-axis shows the goal completion at each step. Solid lines represent results from the $exp_{\text{hint}}$ setting, while dotted lines represent results from the $exp_{\text{base}}$ setting.
@@ 6 Additional Analysis
To scrutinize the performance differences between various models in VisEscape, we conducted additional analysis on action repetition and impact of Memory module.
@@
InternVL2.5-38B Image 3.3 9.02 0.90 292.8 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.70 295.9 35.0 53.76 38.14 7.08 245.3
@@
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.20 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 1.76 0.0 - 15.0 34.79 64.97 1.47 289.8
@@
Table 4: Comparison of scores between different run types: using only MLLM(gets “Image” as input) and using a mixture of MLLM with LLM(gets image “Caption” as input). Metrics improved with caption inputs are in bold. Base LLM refers to the language backbone underlying the corresponding MLLM. Base LLM(R1) for InternVL2.5-38B denotes Deepseek-R1-Distill-Qwen-32B. Experiments without steps all ended in failure due to reaching the maximum turn limit.
@@ 6.1 Repetitive ineffective action patterns
To analyze why higher success rates and goal completion correlate with reduced average steps in Table 3 despite the expectation that more turns would be required for higher SR and GC, we noticed that lower-performing models frequently repeated the same action in identical situations on their trajectories. Figure 7 shows that these redundant attempts are more prevalent in lower-performing models, $exp_{\text{base}}$ and $exp_{\text{hint}}$ both. Consequently, they accumulate a longer trajectory, which in turn leads to lower SPL scores.
Figure 7: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (from 2 to 10+ repetitions) on $exp_{\text{base}}$. The top and bottom graphs represent results from $exp_{\text{hint}}$ and $exp_{\text{base}}$, respectively. The X-axis represents the number of duplicate attempts, and the Y-axis shows the ratio of such attempts.
6.2 Ablation of Memory module
The key characteristic of VisEscape —requiring the agent to retain and utilize information collected throughout the trajectory—highlights memory as essential for effectively navigating partially observable environments. To investigate how and to what extent the Memory module contributes to VisEscaper, we conduct ablations on the Memory module, and analyze the degree of exploration for each condition, with two models-InternVL2.5-34B and LLaVA-v1.6-34B. For ablation, instead of using a Memory module, we enumerated and fed all unique observations that had been observed up to the present to ReAct module.
Table 5 show a significant decrease across metrics due to absence of Memory module. Notably, SR and SPL decreased more significantly than GC, indicating that without the Memory module, the agent struggles to maintain the vast information needed to eventually escape the room.
Model Memory SR(%) GC(%) HR(%) SPL(%) Step
@@ 6.2 Ablation of Memory module
Table 5: Ablation results in $exp_{\text{hint}}$comparing performance in the presence and absence of memory modules.
@@ 6.3 Replacing VLM with LLM
Therefore, we compared the results between this paper’s default setting—where the VLM directly receives visual observations—and the Socratic Models methodology, where the LLM instead receives captions generated by the VLM. For controlled experiments, we used only the LLM that served as backbone language model of corresponding VLM. Results are presented in Table 4.
Among open-source models, InternVL2.5-38B, which demonstrated performance comparable to GPT-4o-mini in the $exp_{\text{hint}}$ setting, showed decrease of nearly half in success rate in both $exp_{\text{base}}$ and $exp_{\text{hint}}$ settings when LLM was adopted instead of VLM, with significant deterioration in all other metrics as well. In contrast, the remaining open-source models demonstrated an average x3.1 increase in SR. A particularly significant observation pertains to the case of InternVL2.5-38B; upon substitution of the LLM from Qwen-32B-Instruct to Deepseek-R1-Distill-Qwen-32B, the identical model architecture but having undergone reasoning capability distillation from state-of-the-art reasoning model (Deepseek-R1 [14]), thus possessing remarkably enhanced reasoning capabilities—there was a marked enhancement in performance metrics across all evaluation criteria. These results suggest that: 1.Many VLMs struggle with complex reasoning in multimodal contexts—a critical capability for VisEscape, and 2.Integrating strong reasoning capabilities of LLMs with visual perception of VLMs can lead to enhanced multimodal reasoning capability.
@@ 7 Conclusion
In this study, we introduced VisEscape, a novel benchmark of escape room scenarios crafted to rigorously test multimodal agents’ abilities to effectively explore, reason, and make decisions in interactive and dynamical environments. We demonstrated that existing multimodal models struggle significantly under these demanding conditions. To address this issue, we proposed VisEscaper framework, achieving substantial improvements in both effectiveness and efficiency. Our findings highlight the critical importance of memory management and adaptive reasoning capacity for multimodal agents, providing pathways toward developing more capable, autonomous AI systems for complex, real-world problem-solving tasks.
@@ 8 Limitation
While VisEscape presents a valuable benchmark for testing exploration-driven problem-solving abilities of multimodal agents, the specific task of escaping rooms may not fully generalize to real-world environments where agents encounter more complex situations and more unfamiliar objects. Additionally, since the action space of VisEscape is discrete rather than continuous, there is a need to convert the conclusions drawn from multiple modules into a feasible continuous action sequence for real-world execution.
@@
  • In appendix B, we provide comprehensive details about our experimental setup, and the prompt template used.
  • In appendix C, we provide additional information about our data construction pipeline.
@@ A.1. Tendency of brute-force approaches
In Figure 8, we show that low-performing agents exhibited meaningless, repetitive action patterns. However, upon analyzing the trajectories of various agents, we found that, when attempting to solve visual quizzes, smaller models frequently resorted to brute-force strategies—randomly trying different passwords or sequences without grounding their actions in any meaningful clues. For example, for problems requiring agents to arrange a given set of numbers in the correct order, some agents repeatedly attempt various possible combinations. To verify whether this behavior constituted brute-force attempts, we directly annotated whether the reasoning process and the corresponding attempted-answer pairs indicated a brute-force approach or not. Specifically, we annotated according to the following rules: (1) if the reasoning was grounded in meaningful clues or observations, then we did not annotate it as brute-force. (2) if the attempted answers involved repeatedly trying different arrangements or combinations without valid justification, then we annotated it as brute-force.
Table 6 demonstrates that while Claude-3.5-Sonnet and GPT-4o typically grounded their puzzle-solving attempts in meaningful visual clues, GPT-4o-mini, InternVL-2.5-38B and LLaVA-v1.6-34B frequently resorted to brute-force approaches. Additionally, models that made greater effort to identify reasoning justifications for their answers exhibited higher success rates in actual experiments. This behavior partially explains why these lower-performing models consistently achieved lower SPL and step efficiency scores—they make more attempts and fail more frequently on quizzes.
Model Brute-force solved ratio(%) Solved quiz ratio(%)
@@
Table 6: Results on puzzle-solving. Brute-force solved ratio denotes the ratio of brute-force attempts among all answers submitted for numeric or alphabetic lock puzzles. Solved quiz ratio denotes the ratio of correctly solved locks to the total number of locks in VisEscape on $exp_{\text{base}}$.
A.2. Visualization of exploration without Memory module
According to Table 5, removing the Memory module notably reduced VisEscaper performance. Beyond performance metrics, we hypothesize that among the three components of Memory module—particularly the exploration memory component, which tracks already explored versus unexplored objects—encourages agents to seek out unseen scenes or objects. To verify this, we conducted an additional ablation experiment in which we specifically removed the exploration memory, while keeping the other two components (structured spatial memory and salient action memory) intact within the Memory module. Figure 9 shows that InternVL-2.5-38B exhibited only a slight difference in exploration of unique scenes, whereas LLaVA-1.6-34B shows a substantial decrease.
@@ A.3. Visualization of interactions in VisEscape
Figure 8 is a visualization of interactions. The size of each word is proportional to its frequency of appearance within the game.
Figure 8: Wordcloud of all defined types of actions in VisEscape. We excluded ”open” and ”close”, which are defined on most receptacles by default.
Figure 9: The proportion of essential scenes observed by the agent at least once within the initial 100 steps. The dotted lines denote models without exploration memory, while solid lines denote VisEscaper.
@@ A.4. Statistics for image captions
Since the Memory module uses image captions to construct and manage long-term memory, it is crucial to evaluate how accurately the VLM captures essential information about the current state of the scene. To evaluate captions, we define the ground-truth (GT) state as the complete set of all receptacles and items present in an image scene obtained from the internal game engine. We then parse the generated captions and examine whether they contain the information matching the GT state. Table 7 shows that captions generated by InternVL2.5-38B provide the most accurate information in the most concise manner, even when compared to GPT-4o.
@@
InternVL2.5-38B 75.1 38.5
Table 7: Accuracy indicates whether the caption accurately describes the current game state. Avg. Length denotes the average length of all captions generated by each model for every unique scene.
@@
Model(MLLM / LLM) Input Modality SR(%) GC(%) SPL(%) Step SR(%) GC(%) HCR(%) SPL(%) Step
@@
InternVL2.5-38B Image 3.3 9.02 0.90 292.8 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.70 295.9 35.0 53.76 38.14 7.08 245.3
@@
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.20 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 1.76 0.0 - 15.0 34.79 64.97 1.47 289.8
@@
Table 8: Full results are shown in Table 4. For GPT-4o and GPT-4o-mini, we used the same model but varied the input modality type, since they do not provide language models that accept text input only.
@@ A.5. Full Results of Table 4
Table 8 shows results including GPT-4o and GPT-4o-mini, which were omitted in Table 4. When GPT-4o-mini used captions instead of images as input, there was a slight improvement across all metrics, but it did not show a significant difference compared to other open-source models. Specifically, GPT-4o’s performance decreased in $exp_{\text{base}}$ but slightly improved in $exp_{\text{hint}}$.

### v2

@@ 5.4 Results
1. Even advanced models cannot achieve success without any hints. Table 3 presents the results of various VLMs integrated within the VisEscaper framework under both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Without any hints ($exp_{\text{base}}$), even Claude-3.5-Sonnet achieves only a 21.7% escape success rate and 32.37% goal completion, which is still 3-4 times lower than the human scores. Considering that humans show a minimal difference between the achievement rate and the final escape success rate (95.80 for GC and 82.5 for SR), the results from all models demonstrate that despite making some progress within each room, they cannot guarantee long-horizontal continuous success when no guidance is provided.
2. High-performing models show a drastic increase in all metrics, achieving objectives with fewer hints and within fewer turns. When hints are provided to stuck agents ($exp_{\text{hint}}$), there is a significant increase in success rate and goal completion across all models. Notably, Claude-3.5-Sonnet and GPT-4o escaped rooms using less than half the average number of steps compared to other models, making them much more efficient at path optimization, which leads to the greatest improvement, especially in SPL, among all metrics. Also, smaller models still improve substantially with hints, but rely on them more frequently, as shown by higher HCR.
3. High-performing models demonstrate accelerated goal completion rates in later stages compared to early stages. Figure 6 illustrates the progression of GC in Table 3 across all rooms during the first 100 steps. Notably, while Claude-3.5-Sonnet performs best overall, it exhibits a distinctly accelerated trajectory after approximately 30-40 steps. This acceleration indicates that strong models better manage accumulated room-state changes, effectively leveraging previous observations to make strategic and efficient decisions later in the trajectory.
Additionally, models exhibiting rapid progress (Claude-3.5-Sonnet and GPT-4o) used fewer hints compared to other models (HCR of $exp_{\text{hint}}$ in Table 3). This indicates that rapid progress by high-performing models does not stem from heavy reliance on hints; instead, they use hints selectively only to resolve specific obstacles and, once past these obstacles, accelerate progress by effectively leveraging information accumulated up to that point.
Figure 6: Average GC (Goal Completion) progression across all rooms and experiments during the first 100 steps. The X-axis represents the number of steps, and the Y-axis represents the goal completion at each step. Solid lines represent results from the $exp_{\text{hint}}$ setting, while dotted lines represent those from the $exp_{\text{base}}$ setting.
@@ 6 Additional Analysis
To scrutinize the performance differences between various models in VisEscape, we conduct additional analysis on action repetition and impact of the Memory module.
@@
InternVL2.5-38B Image 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.67 295.9 35.0 53.76 38.14 7.08 245.3
@@
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
@@
Table 4: Comparison of scores between different run types: using only VLM (which receives an image as input) and using a mixture of VLM with LLM (which receives a caption as input). Metrics improved with caption inputs are highlighted in bold. Base LLM refers to the language backbone underlying the corresponding VLM. Base LLM (R1) for InternVL2.5-38B denotes Deepseek-R1-Distill-Qwen-32B. Experiments without step counts all ended in failure due to reaching the maximum turn limit.
@@ 6.1 Repetitive Ineffective Action Patterns
To analyze why higher success rates and goal completion correlate with reduced average steps in Table 3 despite the expectation that more steps would be required for higher SR and GC, we observe that lower-performing models frequently repeat the same action in identical situations throughout their trajectories. Figure 7 shows that these redundant attempts are more prevalent in lower-performing models, including both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Consequently, they accumulate longer trajectories, which in turn leads to lower SPL scores.
Figure 7: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions) on $exp_{\text{base}}$ setting. The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
[…]
@@
Table 5: Ablation results in $exp_{\text{hint}}$ comparing performance in the presence and absence of memory modules. The scores of all metrics decrease significantly when memory modules is absent.
6.2 Ablation of the Memory Module
The key characteristic of VisEscape —requiring the agent to retain and utilize information collected throughout the trajectory—highlights memory as essential for effectively navigating in partially observable environments. To investigate how and to what extent the Memory module contributes to VisEscaper, we conduct ablations on the Memory module, and analyze the degree of exploration for each condition using two models: InternVL2.5-38B and LLaVA-v1.6-34B. For the ablation, instead of using a Memory module, we enumerate all unique observations made up to the current step and feed them directly into the ReAct module. Table 5 shows a significant decrease across metrics due to absence of the Memory module. Notably, SR and SPL decreased more significantly than GC, indicating that without the Memory module, the agent struggles to retain the extensive information required to eventually escape the room.
@@ 6.3 Replacing VLM with LLM
Therefore, we compared the results between this paper’s default setting—where the VLM directly receives visual observations—and the Socratic Models methodology, where the LLM instead receives captions generated by the VLM. For controlled experiments, we used only the LLM that served as backbone language model of corresponding VLM  [^2] Refer to Appendix B.2. for the names of the corresponding LLMs for each VLM.. Results are presented in Table 4.
Among open-source models, InternVL2.5-38B, which demonstrated performance comparable to GPT-4o-mini in the $exp_{\text{hint}}$ setting, showed a decrease of nearly half in success rate in both $exp_{\text{base}}$ and $exp_{\text{hint}}$ settings when LLM was adopted instead of VLM, with significant deterioration in all other metrics as well. In contrast, the remaining open-source models demonstrated an average x3.1 increase in SR. A particularly significant observation pertains to the case of InternVL2.5-38B; upon substituting the LLM Qwen-32B-Instruct with Deepseek-R1-Distill-Qwen-32B, which shares an identical model architecture but has undergone reasoning capability distillation from a state-of-the-art reasoning model (Deepseek-R1 [15]) and thus possesses remarkably enhanced reasoning capabilities, all evaluation metrics improved notably. These results suggest that: 1. Many VLMs struggle with complex reasoning in multimodal contexts—a critical capability for VisEscape, and 2. Integrating strong reasoning capabilities of LLMs with visual perception of VLMs can lead to enhanced multimodal reasoning capability.
@@ 7 Conclusion
In this study, we introduce VisEscape, a novel benchmark of escape room scenarios crafted to rigorously test multimodal agents’ abilities to effectively explore, reason, and make decisions in interactive and dynamic environments. We demonstrate that existing multimodal models struggle significantly under these demanding conditions. To address this issue, we propose VisEscaper framework, achieving substantial improvements in both effectiveness and efficiency. Our findings highlight the critical importance of memory management and adaptive reasoning capacity for multimodal agents, providing pathways toward developing more capable, autonomous AI systems for complex, real-world problem-solving tasks.
@@ 8 Limitation
While VisEscape presents a valuable benchmark for testing the exploration-driven problem-solving abilities of multimodal agents, the specific task of room escapes may not fully generalize to real-world environments where agents encounter more complex situations and more unfamiliar objects. Additionally, since the action space of VisEscape is discrete rather than continuous, there is a need to convert the conclusions drawn from multiple modules into feasible continuous action sequences for real-world execution.
9 Acknowledgements
This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name: Development of multimodal UX evaluation platform technology for XR spatial responsive content optimization, Project Number: RS-2024-00361757).
@@ 9 Acknowledgements
  • In Appendix B, we provide comprehensive details about our experimental setup and the prompt template used.
  • In Appendix C, we provide additional information about our data construction process.
@@ A.1. Tendency of Brute-Force Approaches
In Figure 7, we show that low-performing agents exhibited meaningless, repetitive action patterns. However, upon analyzing the trajectories of various agents, we found that, when attempting to solve visual quizzes, smaller models frequently resorted to brute-force strategies—randomly trying different passwords or sequences without grounding their actions in any meaningful clues. For example, for problems requiring agents to arrange a given set of numbers in the correct order, some agents repeatedly attempted various possible combinations. To verify whether this behavior constituted brute-force attempts, we directly annotated whether the reasoning process and the corresponding attempted-answer pairs indicated a brute-force approach or not. Specifically, we annotated actions according to the following rules: (1) If the reasoning was grounded in meaningful clues or observations, we did not annotate it as brute-force. (2) If the attempted answers involved repeatedly trying different arrangements or combinations without valid justification, then we annotated them as brute-force.
Table 6 demonstrates that while Claude-3.5-Sonnet and GPT-4o typically grounded their puzzle-solving attempts in meaningful visual clues, GPT-4o-mini, InternVL2.5-38B and LLaVA-v1.6-34B frequently resorted to brute-force approaches. Additionally, models that made a greater effort to identify reasoning justifications for their answers exhibited higher success rates in actual experiments. This behavior partially explains why these lower-performing models consistently achieved lower SPL and step efficiency scores—they made more attempts and fail more frequently on quizzes.
Model Brute-force attempt ratio(%) Solved quiz ratio(%)
@@
Table 6: Results for puzzle-solving. Brute-force attempt ratio denotes the ratio of brute-force attempts among all answers submitted for numeric or alphabetic lock puzzles. Solved quiz ratio denotes the ratio of correctly solved locks to the total number of locks in VisEscape under the $exp_{\text{base}}$ setting.
A.2. Effect of the Exploration Memory
According to Table 5, removing the Memory module notably reduced VisEscaper performance. Beyond performance metrics, we hypothesize that among the three components of Memory module, particularly the Exploration memory component, which tracks already explored versus unexplored scenes or objects, encourages agents to seek out unseen scenes or objects. To verify this, we conducted an additional ablation experiment in which we specifically removed the Exploration memory, while keeping the other two components (Structured spatial memory and Salient action memory) intact within the Memory module. For both settings, we measured the proportion of scenes observed by the agent from the set of essential observations (as determined by the oracle trajectory) required to successfully complete the game. Figure 9 shows that InternVL2.5-38B exhibited only a slight difference in exploration of unique scenes, whereas LLaVA-v1.6-34B shows a substantial decrease.
@@ A.3. Visualization of Interactions in VisEscape
Figure 8 is a visualization of interactions defined in VisEscape. The size of each word is proportional to its frequency of appearance within the game.
Figure 8: Wordcloud of all types of actions defined in VisEscape. We excluded “open” and “close”, which are available on most receptacles.
Figure 9: Proportion of essential scenes observed by the agent at least once within the initial 100 steps. Solid and dotted lines denote results with and without exploration memory, respectively. Removing exploration memory significantly reduces the number of observed essential scenes in both models.
@@ A.4. Statistics for Image Captions
Since the Memory module uses image captions to construct and manage long-term memory, it is crucial to evaluate how accurately the VLM captures essential information about the current state of the scene. To evaluate captions, we define the ground-truth (GT) state as the complete set of all receptacles and items present in an image obtained from the internal game engine. We then parse the generated captions and examine whether they contain the information consistent with the GT state. Table 7 shows that captions generated by InternVL2.5-38B provide the most accurate information in the most concise way, even when compared to GPT-4o.
@@
InternVL2.5-38B 75.1 38.50
Table 7: Evaluation on captions generated by each VLMs. Accuracy indicates whether the caption accurately describes the current game state. Avg. Length denotes the average length of all captions generated by each model for every unique scene.
@@
Model(VLM / LLM) Input Modality SR(%) GC(%) SPL(%) Step SR(%) GC(%) HCR(%) SPL(%) Step
@@
InternVL2.5-38B Image 0.0 6.20 0.0 - 66.7 79.23 41.69 12.23 210.5
InternVL2.5-38B + Base LLM Caption 1.7 9.57 0.67 295.9 35.0 53.76 38.14 7.08 245.3
@@
InternVL2-40B + Base LLM Caption 0.0 3.88 0.0 - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
@@
Table 8: Full results are shown in Table 4. For GPT-4o and GPT-4o-mini, we used the same model but varied the input modality type, since they do not provide language models that receives text input only.
@@ A.5. Full Results of Table 4
Table 8 includes results for GPT-4o and GPT-4o-mini, which were omitted in Table 4. When GPT-4o-mini used captions instead of images as input, there was a slight improvement across all metrics, but it did not show a significant difference compared to other open-source models. Notably, GPT-4o’s performance decreased in $exp_{\text{base}}$ but slightly improved in $exp_{\text{hint}}$ setting.